### v1

@@ Abstract
Fraud in public procurement remains a persistent challenge, especially in large, decentralized systems like Brazil’s Unified Health System. We introduce Heron’s Information Coefficient (HIC), a geometric measure that quantifies how subgraphs deviate from the global structure of a network. Applied to over eight years of Brazilian bidding data for medical supplies, this measure highlights collusive patterns that standard indicators may overlook. Unlike conventional robustness metrics, the Heron coefficient focuses on the interaction between active and inactive subgraphs, revealing structural shifts that may signal coordinated behavior, such as cartel formation. Synthetic experiments support these findings, demonstrating strong detection performance across varying corruption intensities and network sizes. While our results do not replace legal or economic analyses, they offer an effective complementary tool for auditors and policymakers to monitor procurement integrity more effectively. This study demonstrates that simple geometric insight can reveal hidden dynamics in real-world networks better than other Information Theoretic metrics.
@@
We applied the measure to eight years of Brazilian health bidding data (2013–2021), revealing patterns consistent with collusion in COVID-era procurement.
Artificial network tests confirm detection stability when scaled.
Supports auditing by highlighting unusual network motifs and links with greater sensitivity than other metrics.
@@ 1 Introduction
The outbreak of the COVID-19 pandemic in 2020 brought forth an unprecedented demand for medical supplies worldwide, causing potential risks to the availability of essential healthcare items [1, 2, 3]. This surge in demand resulted in fierce competition among suppliers, particularly for individual protection materials like hospital aprons, rubber gloves, and masks. In Brazil, where health is constitutionally guaranteed as a universal right, the government is responsible for procuring and distributing medical supplies for over 210 million citizens through the Brazilian Unified Health System (SUS) [4, 5, 6, 7, 8, 9, 10].
@@
To address these issues, we propose a measure called Heron’s Information Coefficient (HIC) to gain insights into bidding processes for health-related items in Brazil. By understanding network structures and dynamics, we aim to identify potential collusion and enhance the transparency and integrity of public procurement. As such, we consider both the topological connectivity and the degree of protection provided to the system’s elements. Complex network theory provides effective tools to analyze various dynamic systems, enabling a better understanding of transportation, diffusion, synchronization, and social networks, among others [12, 13, 14, 15, 16, 17, 18, 19, 20].
@@
The proposed methodology was tested on both real and artificial networks, revealing insightful behavior in artificial networks and highlighting unusual patterns in the interactions among co-bidding companies for medical supplies in Brazil. The results from this study can provide valuable information to control agencies, aiding their supervision of public procurement processes and combating corruption effectively.
@@ 2 Methods
A network is a set of vertices, $V$, connected by links, $E$. Generally, the network is just the structure over which a time-dependent dynamical process occurs, and its elements must not be available to operate during all periods. In this way, at a given instant $t\geq 0$, the original network $G=(V,E)$ can be decomposed into two components representing the active and inactive subgraphs at that instant: $G^{a}(t)=(V^{a}(t),E^{a}(t))$ and $G^{i}(t)=(V^{i}(t),E^{i}(t))$.
@@
If $G^{a}=G$, the whole network is active but very vulnerable to failures as any activity may fail. The opposite, $G^{i}=G$, represents that the whole network is protected, but no dynamics occur over the system. All other intermediate possibilities result in networks that may be more or less similar to the original one. To compare the active and inactive parts and the original network, we propose using the $D$ measure [21], a pseudo-metric between networks. This highly precise network dissimilarity measure compares, through the Jensen-Shannon divergence [22], topological differences between connectivity existing patterns between networks capturing differences between flows of information through shortest paths.
Besides its precision, the $D$ measure also maintains the validity of the triangular inequality, so that, for any given three networks $G_{1}$, $G_{2}$ and $G_{3}$, $D(G_{1},G_{2})+D(G_{2},G_{3})\leq D(G_{1},G_{3})$. Therefore, if at least one of these distances is positive, there are infinite different triangles possessing a perimeter ($2P$) equals three times the average value between their distances: $2P=D(G_{1},G_{2})+D(G_{2},G_{3})+D(G_{1},G_{3})$.
@@
The most symmetrical situation occurs when the distance values are the same, resulting in an equilateral triangle. By extension, all other configurations generate more heterogeneous patterns of distances. Since the largest triangle for a given perimeter is equilateral and the Heron’s formula gives the area of a triangle with sides $D(G_{1},G_{2})$, $D(G_{1},G_{3})$ e $D(G_{2},G_{3})$ as $\sqrt{P(P-D(G_{1},G_{2}))(P-D(G_{1},G_{3}))(P-D(G_{2},G_{3}))}$, we define Heron’s Information Coefficient as the fraction between the area of that triangle and the area of the equilateral triangle with perimeter $2P$:
$\mathcal{H}_{e}(G_{1},G_{2},G_{3})=\frac{3\sqrt{3P\bigl(P-D(G_{1},G_{2})\bigr)\bigl(P-D(G_{1},G_{3})\bigr)\bigl(P-D(G_{2},G_{3})\bigr)}}{P^{2}}$ (1)
@@
Figure 2 shows a multidimensional scaling map of the distances $G_{1}$, $G_{2}$, $G_{3}$ and $G_{4}$ between four small networks. In this representation, we can see that the largest possible triangle formed by triples of these networks has vertices $G_{1}$, $G_{3}$ and $G_{4}$ with similar sides. When combining $G_{1}$, $G_{3}$ and $G_{4}$ no triangle is formed because, in fact, $D(G_{1},G_{3})=D(G_{1},G_{2})+D(G_{2},G_{3})$. Thus, among all the non-isomorphic networks of size four ${\cal H}_{e}(G_{1},G_{3},G_{4})=0.996$ represents its highest possible value while ${\cal H}_{e}(G_{1},G_{2},G_{3})=0$ represents its minimum.
@@ 3 Characterizing Heron’s Information Coefficient’s Behavior
For small $p$ values, the inactive network closely resembles the complete graph, resulting in low HIC values. As $p$ increases, the topologies of the original, active, and inactive networks diverge until a maximum, reached when approximately 20% of the links are active. Beyond this point, HIC gradually decreases until a local minimum in $p=0.5$. At this local minimum, the active and inactive networks are topologically similar subgraphs of the original network, as the probability of a link existence is 50% in both.
@@
The random graph model is a good example but does not tell the whole story. In such a network, each link has an equal probability of being active, a pattern not often presented in the real world. In social networks, for example, some ties tend to be stronger than others. To introduce such heterogeneity while maintaining the same expected number of links, we assign link activation probabilities proportional to their edge betweenness centrality ($b_{e}$), a measure reflecting the number of shortest paths that traverse each edge. Specifically, for a given parameter $\gamma$, the activation probability of an edge is defined as $p_{e}\propto b_{e}^{\gamma}$.
Figure 4(A) shows an example of an active–inactive configuration in a network with 100 vertices, where thicker and redder edges denote higher betweenness centrality values. As depicted in Figure 4(B), we computed the average HIC across 100 random seeds for different $\gamma$ values. Negative $\gamma$ values correspond to preferential activation of low-betweenness edges, whereas positive $\gamma$ values emphasize highly central ones. The results reveal that HIC tends to increase when edges with higher betweenness are more likely to be active, reflecting more efficient structural differentiation between active and inactive subgraphs.
@@
Overall, a high Heron coefficient is associated with the system’s efficiency in transferring information. We observe high values in systems whose active and inactive parts are equidistant from the original network and from each other, which indicates different connectivity patterns without losing information from the original network. In the network represented in Figure 4, the preferential attachment to high edge betweenness edges ensures that the information transfer during the activity period travels the network more efficiently in different points, whereas the inactivity patterns protect the network’s periphery. Simultaneously, however, important links in the system - those with high betweenness centrality, for instance - may fail. In this case, a possible strategy would be to interchange the active and inactive networks while retaining Heron’s information coefficient so that the information that reaches the system’s periphery rapidly through the most central active network can further spread to peripheral nodes.
@@
To this end, we conducted experiments across multiple network topologies. The scale-free Barabási–Albert model grows through preferential attachment, producing hubs with high degree centrality and a topology resilient to random failures but highly vulnerable to targeted attacks on hubs. The small-world or Watts–Strogatz model combines local clustering with short global path lengths, which confer it balanced robustness across attack strategies. On the other hand, the Erdős-Rényi network, whose robustness depends mainly on edge density, served as a baseline due to its rapid fragmentation under targeted attacks. We also analyzed a community-structured network, characterized by densely connected clusters with sparse inter-community links, to evaluate the effects of bridge removal on global connectivity. In contrast, the hub-and-spoke topology is a centralized structure dominated by a single hub, extremely fragile under targeted removal of that central node, emphasizing dependence on a single point of failure. Finally, the real bidding network exemplifies the problem we aim to tackle, capturing genuine structural patterns observed in procurement systems.
@@
Figure 5: Sensitivity comparison of the edge removal across relevant metrics (HIC, degree robustness, betweeness robustness, clustering, and modularity) for different network types (scale-free, small-world, Erdős–Rényi, community-structured, hub-and-spoke, and a real bidding network. The real network corresponds to the 2016 Surgical Masks network, where our metric outperformed the others, as it consistently did for Brazilian bidding networks.
@@
Each network underwent controlled node removal experiments, in which 10% of the nodes were sequentially removed in descending order of degree. Across all network types, our coefficient consistently exhibited higher sensitivity to structural degradation than the other evaluated metrics. Our coefficient’s performance reinforces its suitability for detecting covert or collusive groups, surpassing state-of-the-art methods based on edge betweenness in capturing critical structural vulnerabilities.
@@
Extensive testing across multiple network configurations confirmed the expected behavior and robustness of our coefficient. Together, these findings demonstrate that Heron’s Information Coefficient captures both the structural balance between active and inactive subgraphs and the dynamic efficiency of information flow within the system. Its sensitivity to perturbations across a variety of topologies highlights our coefficient’s potential.

### v2

@@ Abstract
Fraud in public procurement remains a persistent challenge, especially in large, decentralized systems like Brazil’s Unified Health System. We introduce Heron’s Information Coefficient (HIC), a geometric measure that quantifies how subgraphs deviate from the global structure of a network. Applied to over eight years of Brazilian bidding data for medical supplies, this measure highlights possible collusive patterns that standard indicators may overlook. Unlike conventional robustness metrics, the Heron coefficient focuses on the interaction between active and inactive subgraphs, revealing structural shifts that may signal coordinated behavior, such as cartel formation. Synthetic experiments support these findings, demonstrating strong detection performance across varying corruption intensities and network sizes. While our results do not replace legal or economic analyses, they offer an effective complementary tool for auditors and policymakers to monitor procurement integrity more effectively. This study demonstrates that simple geometric insight can reveal hidden dynamics in real-world networks better than other Information Theoretic metrics.
@@
We applied the measure to eight years of Brazilian health bidding data (2013–2021), revealing similar abnormal patterns before, during and after COVID-era procurement, possible indicators of collusion.
Artificial network tests confirm anomaly detection stability when scaled.
HIC supports auditing by highlighting unusual network motifs and links with greater sensitivity than other metrics.
@@ 1 Introduction
The outbreak of the COVID-19 pandemic in 2020 brought forth an unprecedented demand for medical supplies worldwide, causing potential risks to the availability of essential healthcare items [37, 35, 28]. This surge in demand resulted in fierce competition among suppliers, particularly for individual protection materials such as hospital aprons, rubber gloves, and masks. In Brazil, where health is constitutionally guaranteed as a universal right, the government is responsible for the procurement and distribution medical supplies for over 210 million citizens through the Brazilian Unified Health System (SUS) [30, 31, 18, 39, 11, 5, 13].
@@
To address these issues, we propose a measure called Heron’s Information Coefficient (HIC) to gain insights into bidding processes for health-related items in Brazil. By understanding network structures and dynamics, we aim to identify potential collusion and enhance the transparency and integrity of public procurement. As such, we consider both the topological connectivity and the degree of protection provided to the system’s elements. Complex network theory provides effective tools for analyzing various dynamic systems, enabling a better understanding of transportation, diffusion, synchronization, and social networks, among others [4, 6, 7, 29, 34, 36, 41, 43, 44].
@@
The proposed methodology was tested on both real and artificial networks, revealing insightful behavior in artificial networks and highlighting unusual patterns in the interactions among co-bidding companies for medical supplies in Brazil. The results of this study can provide valuable information to control agencies, aiding their supervision of public procurement processes and combating corruption effectively.
@@ 2 Methods
A network is a set of vertices, $V$, connected by links, $E$. Generally, the network is just the structure over which a time-dependent dynamical process occurs, and its elements must not be available to operate during all periods. In this way, at a given instant $t\geq 0$, the original network $G=(V,E)$ can be decomposed into two components that represent the active and inactive subgraphs at that instant: $G^{a}(t)=(V^{a}(t),E^{a}(t))$ and $G^{i}(t)=(V^{i}(t),E^{i}(t))$.
@@
If $G^{a}=G$, the entire network is active but very vulnerable to failures as any activity can fail. Conversely, the condition $G^{i}=G$ implies a system in topological stasis, fully protected yet dynamically inert. All other intermediate possibilities result in networks that may be more or less similar to the original one. To compare the active and inactive parts and the original network, we employ the $D$ measure [32], a pseudo-metric between networks. This highly precise measure of network dissimilarity relies on the Jensen-Shannon divergence [25] to assess topological differences in connectivity patterns, effectively capturing the divergence in information flow along shortest paths.
Formally, let $P_{i}=\{p_{i}(j)\}$ denote the distance distribution of node $i$, where $p_{i}(j)$ represents the fraction of nodes located at a shortest-path distance $j$. We quantify the topological heterogeneity of the system via the Network Node Dispersion (NND), defined as $\mathrm{NND}(G)=\mathcal{J}(P_{1},\ldots,P_{N})/\log(d+1)$, where $\mathcal{J}$ is the Jensen-Shannon divergence and $d$ is the network diameter. The $D$-measure synthesizes three distinct topological descriptors: (i) the global difference between the averaged node-distance distributions, $\mu_{G}$ and $\mu_{G^{\prime}}$; (ii) the aforementioned NND, which captures local variance in distance profiles; and (iii) a spectral comparison of the $\alpha$-centrality values of the graphs and their complements. This final term, also computed via Jensen-Shannon divergence, is critical for resolving isomorphisms in distance-regular structures and distinguishing between complete graphs of varying sizes or systems containing disconnected components. As such, the distance $D(G,G_{0})$ between two networks enables a comparison based on global distance structure rather than local connectivity alone:
[…]
where $N$ and $M$ are the sizes of $G$ and $G^{\prime}$, respectively, and $G^{c}$ indicates the complement of $G$. The weights $w_{1}$, $w_{2}$ and $w_{3}$ satisfy $w_{1}+w_{2}+w_{3}=1$, and the authors of the D-measure [32] show that the choice of the weights does not change the metric character and present a detailed discussion on the weight selection. Typically, $w_{3}$ is set to 0.1, as a high value for this term would distort the D values, minimizing the power of the NND in detecting these important topological changes. Due to the computational complexity added with this term, $w_{3}$ can can be avoided without losing precision - unlike $w_{1}$ and $w_{2}$, which are considered equally important and thus set to $w_{1}=w_{2}=0.5$. By setting $w_{3}=0$, we rely entirely on terms $w_{1}$ and $w_{2}$, which are much better at detecting massive structural breaks, such as phase transitions or disconnections.
In addition to its precision, the $D$-measure also satisfies the triangular inequality, so that, for any given three networks $G_{1}$, $G_{2}$ and $G_{3}$, $D(G_{1},G_{3})\leq D(G_{1},G_{2})+D(G_{2},G_{3})$. By satisfying positive definiteness, symmetry, and the triangular inequality, the $D$-measure effectively projects the complex, non-Euclidean topology of graph structures into a searchable metric space [23]. This projection ensures that the familiar laws of geometry and algebra remain valid, allowing us to reason about the distance between complex systems as rigorously as we reason about distance in the physical world.
The operational impact of this geometric validity cannot be overstated. The triangular inequality is the cornerstone of efficient high-dimensional search, since it enables both the computational scalability and the semantic definition of structural anomalies. It allows us to bound the distance between two objects without explicitly comparing them: if $G_{1}$ is close to $G_{2}$, and $G_{2}$ is far from $G_{3}$, the inequality mathematically guarantees that $G_{1}$ is also far from $G_{3}$. Without the pruning power granted by this axiom, the index structures required to identify outliers in billion-node graphs would be computationally intractable. Thus, the inequality transforms the concept of an ”anomaly” from a subjective observation into a geometric fact: an object that physically cannot be embedded within the dense manifold of the distribution without violating the axioms of the space.
@@
Given the triangular inequality, if at least one of these distances is positive, there are infinite different triangles possessing a perimeter ($2P$) equal to three times the average value between their distances: $2P=D(G_{1},G_{2})+D(G_{2},G_{3})+D(G_{1},G_{3})$. The most symmetrical situation occurs when the distance values are the same, resulting in an equilateral triangle in the metric space. By extension, all other configurations generate more heterogeneous distance patterns and the perimeter of their triangles becomes a measure of the total structural divergence within the triplet. Since the largest triangle for a given perimeter is equilateral and the Heron’s formula gives the area of a triangle with sides $D(G_{1},G_{2})$, $D(G_{1},G_{3})$ and $D(G_{2},G_{3})$ as $\sqrt{P(P-D(G_{1},G_{2}))(P-D(G_{1},G_{3}))(P-D(G_{2},G_{3}))}$, we define Heron’s Information Coefficient as the fraction between the area of that triangle and the area of the equilateral triangle with perimeter $2P$:
$\mathcal{H}_{e}(G_{1},G_{2},G_{3})=\frac{3\sqrt{3P\bigl(P-D(G_{1},G_{2})\bigr)\bigl(P-D(G_{1},G_{3})\bigr)\bigl(P-D(G_{2},G_{3})\bigr)}}{P^{2}}.$ (2)
@@
Figure 2 shows a multidimensional scaling map of the distances $G_{1}$, $G_{2}$, $G_{3}$ and $G_{4}$ between four small networks. In this representation, we can see that the largest possible triangle formed by triples of these networks has vertices $G_{1}$, $G_{3}$ and $G_{4}$ with similar sides. When combining $G_{1}$, $G_{3}$ and $G_{4}$ no triangle is formed because, in fact, $D(G_{1},G_{3})=D(G_{1},G_{2})+D(G_{2},G_{3})$. Thus, among all non-isomorphic networks of size four, ${\cal H}_{e}(G_{1},G_{3},G_{4})=0.996$ represents its highest possible value while ${\cal H}_{e}(G_{1},G_{2},G_{3})=0$ represents its minimum.
@@ 3 Characterizing Heron’s Information Coefficient’s Behavior
For small $p$ values, the inactive network closely resembles the complete graph, resulting in low HIC values. As $p$ increases, the topologies of the original, active, and inactive networks diverge until a maximum, reached when approximately 20% of the links are active. Beyond this point, the HIC gradually decreases until a local minimum in $p=0.5$. At this local minimum, the active and inactive networks are topologically similar subgraphs of the original network, as the probability of the existence of a link is 50% in both.
@@
The random graph model is a good example, but it does not tell the whole story. In such a network, each link has an equal probability of being active, a pattern that is not often presented in the real world. In social networks, for example, some ties tend to be stronger than others. To introduce such heterogeneity while maintaining the same expected number of links, we assign link activation probabilities proportional to their edge betweenness centrality ($b_{e}$), a measure reflecting the number of shortest paths that traverse each edge. Specifically, for a given parameter $\gamma$, the activation probability of an edge is defined as $p_{e}\propto b_{e}^{\gamma}$.
Figure 4(A) shows an example of an active–inactive configuration in a network with 100 vertices, where the thicker and redder edges denote higher betweenness centrality values. As depicted in Figure 4(B), we computed the average HIC across 100 random seeds for different $\gamma$ values. Negative $\gamma$ values correspond to preferential activation of low-betweenness edges, whereas positive $\gamma$ values emphasize highly central ones. The results reveal that HIC tends to increase when edges with higher betweenness are more likely to be active, reflecting a more efficient structural differentiation between active and inactive subgraphs.
@@
Overall, a high Heron coefficient is associated with the system’s efficiency in transferring information. We observe high values in systems whose active and inactive parts are equidistant from the original network and from each other, which indicates different connectivity patterns without losing information from the original network. In the network represented in Figure 4, the preferential attachment to high edge betweenness edges ensures that the information transfer during the activity period travels the network more efficiently in different points, whereas inactivity patterns protect the network’s periphery. Simultaneously, however, important links in the system - those with high betweenness centrality, for instance - may fail. In this case, a possible strategy would be to interchange the active and inactive networks while retaining Heron’s information coefficient so that the information that reaches the system’s periphery rapidly through the most central active network can further spread to peripheral nodes.
@@
To this end, we conducted experiments across multiple network topologies. The scale-free Barabási–Albert model grows through preferential attachment, producing hubs with high degree centrality and a topology resilient to random failures but highly vulnerable to targeted attacks on hubs. The small-world or Watts–Strogatz model combines local clustering with short global path lengths, which confer it balanced robustness across attack strategies. On the other hand, the Erdős-Rényi network, whose robustness depends mainly on edge density, served as a baseline due to its rapid fragmentation under targeted attacks. We also analyzed a community-structured network, characterized by densely connected clusters with sparse inter-community links, to evaluate the effects of bridge removal on global connectivity. In contrast, the hub-and-spoke topology is a centralized structure dominated by a single hub, extremely fragile under targeted removal of that central node, emphasizing dependence on a single point of failure. Finally, the real bidding network exemplifies the problem we aim to tackle, capturing the genuine structural patterns observed in procurement systems.
@@
Figure 5: Sensitivity comparison of the edge removal across relevant metrics (HIC, degree robustness, betweeness robustness, clustering, and modularity) for different network types (scale-free, small-world, Erdős–Rényi, community-structured, hub-and-spoke, and a real bidding network). The real network corresponds to the 2016 Surgical Masks network, where our metric outperformed the others, as it consistently did for Brazilian bidding networks.
@@
Each network underwent controlled node removal experiments, in which 10% of the nodes were sequentially removed in descending order of degree. Across all types of network, our coefficient consistently exhibited a higher sensitivity to structural degradation than the other metrics evaluated. Our coefficient’s performance reinforces its suitability for detecting covert or collusive groups, surpassing state-of-the-art methods based on edge betweenness in capturing critical structural vulnerabilities.
@@
Extensive testing across multiple network configurations confirmed the expected behavior and robustness of our coefficient. Together, these findings demonstrate that Heron’s Information Coefficient captures both the structural balance between active and inactive subgraphs and the dynamic efficiency of the information flow within the system. Its sensitivity to perturbations across a variety of topologies highlights our coefficient’s potential.
3.1 Heron’s Information Coefficient in collusion detection
Within the broader anomaly detection literature, detecting anomalies in data serves as a critical filter for identifying deviations in high-dimensional systems, spanning domains from financial forensics to social network security. While applicable to various data modalities, we prioritize graph-based approaches due to their unique capacity to model long-range correlations and structural interdependencies. In the context of fraud, for instance, anomalies are rarely isolated events; they are inherently relational. For instance, fraudulent activity typically manifests through specific topological patterns: either as opportunistic contagion, where fraud propagates through established social links; or as organized collusion, where dense subgraphs of actors synchronize their behavior [3].
Historically, detection frameworks relied on heuristic feature engineering and domain-specific statistical models. This paradigm is labor-intensive but also brittle, struggling to generalize to previously unseen anomaly types [26]. While traditional machine learning techniques, such as Support Vector Machines (SVMs) and matrix factorization, attempted to automate feature extraction, they frequently succumb to the curse of dimensionality and prohibitive computational costs when scaling to massive adjacency matrices. Consequently, the field has pivoted toward Deep Learning to capture the non-linear complexity of graph data and learn expressive latent representations capable of separating normal from anomalous manifolds. However, this gain in predictive performance introduces a new critical limitation: these models operate as ”black boxes,” offering superior accuracy but lacking the interpretability required for high-stakes decision-making.
More specifically, the study of covert networks is challenging due to a myriad of reasons, one of which is data incompleteness and unreliability. Publicly available and reliable datasets, such as the Sicilian Mafia dataset [9], are rare and heavily contingent on judicial disclosures. Criminal network research is also uniquely sensitive to sampling biases. For example, random sampling can severely distort the understanding of a network clustered around a few focal nodes. In fact, the collection of complete data is a virtually impossible task, failing to properly capture the complete dynamics of the networks, with bias being inevitable [16]. This uncertainty is compounded by the boundary-specification problem: in the absence of explicit membership rosters, defining the limits of a clandestine network requires inferring latent connections within a secretive, non-cooperative environment.
Moreover, the current detection methods used by law enforcement agencies are restricted from the greater public. Understandably so. Knowing the exact criteria used to identify potential criminal agents could ease organized crime’s efforts in evading capture, in an adversarial adaptation. In the criminology literature, a recurrent approach consists of Network Science techniques. Its primary utility lies in its ability to map dynamic covert systems without relying on a priori organizational templates. Rather than forcing data into presumed hierarchies, graph-theoretical approaches allow the structural properties of the network to emerge from the empirical data.
However, the interpretation of these emergent topologies requires significant nuance to avoid misleading inferences. For instance, the high degree connectivity of an element in the network does not necessarily equate to high status and, conversely, peripheral actors, with low connectivity, may exercise disproportionate control through strategic compartmentalization. For instance, in human trafficking rings, ”madams” may appear peripheral because they operate independently, despite driving the market demand and controlling ”managers” who appear more central in terms of degree centrality [8]. Therefore, structural measures must be contextually grounded. Researchers must must integrate edge attributes and qualitative metadata, recognizing that the absence of a tie carries as much information-theoretic weight as its presence.
Notably, sequential node removal based on betweenness centrality is among the most effective measures for identifying strategic positions within covert or collusive structures [9, 16]. Betweeness produced a greater impact in dismantling a covert network, showing its capability to properly identify relevant elements. This evidence underscores that in covert networks, the capacity to bridge structural holes and control information flow is a more accurate predictor of systemic importance than mere connectivity.
For bidding processes, the collusion between companies is hard to identify and can take multiple forms, using the techniques mentioned in this section [17]. For instance, the recurrence of bids for a particular item [35, 38] can contribute to cartel companies’ prior planning and their subsequent organization in rounds and punishment of the players who fail to cooperate in the next bid. Companies that act in collusion usually participate in a large amount of bidding to share the average earnings or confer an impression of honesty throughout the process [38]. Therefore, we model each bidding item as a network in which each node is a company participating in the bidding process. Each link exists if two companies are competitors in the same contest. To that end, multiple links represent multiple instances where the companies were co-bidding. As such, the bidding process possesses a complex topology structure that is very informative but often difficult to analyze.
The theoretical benchmark of a public bidding network consists of nodes uniformly connected and characterized by the nonexistence of strong and connected communities, leading to a challenging trade-off: extracting connections from this set and retaining those that represent a high-level interaction between companies without overemphasizing what could be simple statistical deviations. In this scenario, Heron’s Information Coefficient is useful due to its quantification of structural imbalance in dynamic networks, revealing signs of the coordinated behavior typical of collusion.
Here, we use the disparity filter algorithm [33] alongside the maximization of HIC to highlight the central players in the BID network. Figure 6 shows how our method works iteratively.
Figure 6: Diagram of our HIC method for collusion detection with a disparity filter. The HIC loop consists of finding the optimal separation threshold $\alpha_{T}$, defined as the point where the structural divergence between the original network ($G$), the backbone ($G_{P}^{\alpha}$), and its complement ($G_{H}^{\alpha}$) is maximized. Geometrically, this corresponds to maximizing the area of the triangle formed by these three states in the D-measure metric space in each iteration. HIC is maximized for a set number of iterations or until the graph is entirely disconnected, depending on the experiment.
The idea of the disparity filter is to associate with each network’s link a weight $\alpha_{ij}$ that represents a significant level provided by the distribution of the local uniform randomness null hypothesis [33]. Relevant links are chosen so that $\alpha_{ij}<\alpha$, $\alpha\in[0,1]$ being a fixed significance level. Instead of arbitrarily choosing the significance levels, we maximize HIC when $\alpha$ varies. Disconnected companies are excluded from the resulting network. The process is repeated until reaching a significance level as small as desired. The selection of the most relevant group is made iteratively, starting with the original network with a significance level equal to 1. A new subgraph is obtained at each step by removing edges that possess a higher significant level. The goal is to find a lower significance level for each step, resulting in a higher HIC. By definition, this happens when the topologies of the active, inactive, and original networks are close to equidistant.
Each company that participated in the bidding process is a node connected by links weighted by the number of times they have competed in the same bidding process. The way nodes interact and connect provides essential information about the system [2]. In particular, information about their connectivity patterns represents a valuable tool on how different dynamics occur in the system. For example, the speed at which information travels within a cohesive community is much greater than in a more sparse one. As such, let $G$ be the original weighted network after applying the disparity filter, for a fixed ${\alpha}\in[0,1]$, and $G_{a}^{\alpha}$ be the network obtained from $G$ by removing links whose weight is greater than or equal to $\alpha$. As such, $G_{i}^{\alpha}=G-G_{a}^{\alpha}$ is the complement of $G_{a}^{\alpha}$ to $G$. We define $\alpha_{T}$ as the smallest value of $\alpha$ such that ${\cal H}_{e}(G,G_{a}^{\alpha},G_{i}^{\alpha})$ is as large as possible. This optimal threshold $\alpha_{T}$ thus identifies the most informative backbone of the network, balancing cohesion and separation between active and inactive subgraphs. In doing so, it reveals the most representative organization of competitive interactions among companies.
To characterize HIC’s behavior in real procurement data, we use a Brazilian dataset of infrastructure projects procurement [17], which documents the procurement activities of the state-controlled oil company Petrobrás between 2002 and 2013. This era encompasses the systemic market manipulation uncovered by the ”Operation Car Wash” investigation, a landmark inquiry that exposed the ”Club of 16” - a cartel of Brazil’s largest construction firms that orchestrated bid-rigging and price - fixing across major infrastructure projects. Despite lacking the temporal aspect, this dataset provides valuable insights from known collusive agents in Brazilian biddings. The original authors established, between machine leaning methods, the superior efficacy of ensemble methods, specifically identifying Extra Trees, Random Forest, and AdaBoost as the top-performing algorithms, with detection rates between 81% and 95%. We reproduced their experiments and tested disparity filters with HIC and degree robustness, given its strong performance in the experiments shown in Figure 5.
The network comprises 272 corporate nodes, of which 47 (17.28%) are identified as corrupt; notably, only 80 companies successfully won at least one of the 101 analyzed auctions. Within the total volume of 683 bids, 128 were flagged as collusive. A defining structural characteristic of this network is its low collusion assortativity ($r=0.1480$). This suggests a deliberate ”cover bidding” strategy, where collusive agents dilute their presence among honest participants to mask topological clustering. Despite this challenging scenario, HIC was able to isolate only collusive agents after 10 iterations, as shown in Figure 7. The only moment our method failed to identify at least 5 collusive agents in its top 10 was in the tenth iteration, when the network only had 4 companies - all of which were corrupt.
Figure 7: Evolution of network composition during iterative backbone extraction for Petrobrás procurement from 2002 to 2013. (Left) Hub Dominance: The count of collusive agents within the top 10 most central nodes identified with HIC, across iterations. The algorithm consistently retains collusive ”ringleaders” in the top ranking, peaking at 100% dominance by Iteration 7. (Right) Network Composition: The stacked bars show the absolute number of nodes remaining at each step, colored by ground truth (Blue = Honest, Red = Collusive). The dashed line tracks the percentage of collusive nodes in the remaining subgraph. As the algorithm iterates, it disproportionately sheds honest nodes (noise), resulting in a monotonic increase in cartel density, reaching 100% purity at Iteration 10.
To validate these findings, we conducted a rigorous sensitivity analysis using Monte Carlo simulations with bootstrap sampling for stability and null models for significance testing. The results reveal a sharp divergence in metric utility. Standard topological measures, specifically Degree Robustness and Modularity, failed to reject the null hypothesis, suggesting they are blind to the cartel’s non-modular structure. In contrast, HIC identified a significant structural anomaly (Z=4.32,$p<0.001$). The observed HIC value ($0.856$) lies far outside the null distribution mean ($\mu\approx 0.69$), confirming that the collusion manifests not as simple clustering, but as a subtle ”information backbone.” This provides compelling statistical evidence that HIC is a superior metric for detecting the subtle coordination characteristic of collusion, which standard topological measures overlook.